### v1

# Look at the Text: Instruction-Tuned Language Models are More Robust Multiple Choice Selectors than You Think

Abstract
Multiple choice questions (MCQs) are commonly used to evaluate the capabilities of large language models (LLMs). One common way to evaluate the model response is to rank the candidate answers based on the log probability of the first token prediction. An alternative way is to examine the text output. Prior work has shown that first token probabilities lack robustness to changes in MCQ phrasing, and that first token probabilities do not match text answers for instruction-tuned models. Therefore, in this paper, we investigate the robustness of text answers. We show that the text answers are more robust to question perturbations than the first token probabilities, when the first token answers mismatch the text answers. The difference in robustness increases as the mismatch rate becomes greater. As the mismatch reaches over 50%, the text answer is more robust to option order changes than the debiased first token probabilities using state-of-the-art debiasing methods such as PriDe. Our findings provide further evidence for the benefits of text answer evaluation over first token probability evaluation.
1 Introduction
The open-ended nature of autoregressive language generation complicates the evaluation of Large Language Models (LLMs). One popular solution to this problem is to prompt LLMs with Multiple Choice Questions (MCQs), which limit the answer space to a few candidate options, thus enabling evaluation by comparing model choices against gold labels. For MCQs, there are two main ways of extracting model choices from their generated text answers: 1) In the text-based approach, the choice is automatically extracted from the text answer either by pattern matching (Wang et al., 2022) or by prompting a strong LLM (Chiang et al., 2023; Li et al., 2023) such as GPT4 (OpenAI et al., 2024). However, pattern matching can often be inaccurate because each model has different response styles, requiring manual feature engineering. This makes pattern matching infeasible when evaluating many tasks and models at the same time. Using a strong proprietary model like GPT-4 as an evaluator, on the other hand, is more flexible but lacks transparency and reproducibility, and also creates high financial costs when we running large-scale evaluations. 2) In comparison, the probability-based approach simplifies the evaluation by ranking the log probabilities assigned to the option IDs (e.g. A/B/C/D) from the first token prediction of the model. This approach is widely adopted in many different benchmarks and LLM evaluation studies (Hendrycks et al., 2020; bench authors, 2023; Liang et al., 2022; Santurkar et al., 2023). However, recent works point to a mismatch between the text- and probability-based approaches (Lyu et al., 2024; Wang et al., 2024), showing that first token probabilities do not match text answers given by models, especially for models fine-tuned on conversational or safety data.
Moreover, recent studies have shown that the probability-based approach lacks robustness – that it is sensitive to linguistic properties of the MCQ prompt (Leidinger et al., 2023) or perturbations such as typos, adding options, word swapping (Tjuatja et al., 2023) and option position (also called selection bias, Dominguez-Olmedo et al., 2023; Zheng et al., 2023).
In this paper, we investigate the robustness of the text-based approach. We compare it with the first token answer as well as the debiased first token answer using the first token debiasing method PriDe (Zheng et al., 2023), by measuring the answer robustness under various prompt perturbations. To avoid the labor of feature engineering and the high cost of using a proprietary model as an evaluator, we fine-tune a Mistral-7b-Instruct model and show that it can reliably and accurately detect the choice from the text answers across models and datasets.  [^1] We will share the dataset and trained classifiers at https://github.com/mainlp/MCQ-Robustness.
By comparing the choice made in the text answer and the first token probability, we show that: (1) The text answer shows small selection bias (§3.3) and high robustness to various sentence perturbations (§3.5) across all the models we examined; (2) The robustness discrepancy between the first token probability and the text answer increases as the mismatch rate between them increases (§3.3, §3.5); (3) When the mismatch rate is high (over 50%), the text answer shows a smaller selection bias than the state-of-art first token debiasing method PriDe (§3.3). As a whole, our work provides further evidence for the benefits of text-based over probability-based MCQ evaluations of LLMs.
2 Mismatch between the probability and the text-based MCQ evaluation
Before the rise of instruction-tuned large language models, it was appropriate to check the probabilities assigned to the options in MCQs by the models for several reasons: 1) Discriminative models such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) are forced to only give probabilities to the available opinions by finetuning on the task with a task-specific classification head; 2) Sequence-to-sequence models such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020) are good at giving the option ID in the next-word prediction since they were specifically trained to perform NLP tasks of various formats but they are not good at “following a user’s intent” in real-life scenarios (Ouyang et al., 2022); 3) Foundation models such as GPT3 (Brown et al., 2020), OPT (Zhang et al., 2022) have poor instruction-following ability which makes the text completion uninterpretable, thus, the prediction of a model is measured by ranking the log probability of the available options.
Techniques like Instruction-Tuning (Wei et al., 2021) and Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) improve the model’s ability to follow the instructions and give helpful responses. Therefore, the model can directly answer the MCQs and give its choice in the text answer. This makes it less appropriate to use probability-based evaluation since the first token probability doesn’t represent the text answer, even specifically asking it to start with the option ID in the prompt (Wang et al., 2024). Figure 1 shows an example of the mismatch between the first token probabilities and the text answer given by the Llama2-7b-Chat model. This mismatch leads to the different performance of a model on a benchmark depending on the evaluation method, as shown in Table 1.
Figure 1: An example mismatch case between the first token probabilities and the text answer given by the Llama2-7b-Chat model.
Model (0-shot) First Token Text Answer Gemma-7b-Inst 30.2 50.8 Llama2-7b-Chat 34.9 43.1 Llama2-13b-Chat 40.2 47.6 Mistral-7b-Inst-0.2 53.2 53.6
Table 1: Performance of the models on MMLU with different evaluation methods: First token probability and Text answer. Text answer achieves better performance in general.
3 Robustness evaluation of the probability and text-based approaches
Several works (Dominguez-Olmedo et al., 2023; Zheng et al., 2023) have shown that large language models have selection bias such as preferring “option A”, therefore they are not robust under position change of the options. To mitigate this bias, PriDe (Zheng et al., 2023) was proposed as a debiasing method to improve the robustness of the model by disentangling the token bias from the first token probability. Tjuatja et al. (2023) shows that LLMs are also sensitive to perturbation in the question such as typos, word swapping and adding more options. However, this line of work only focuses on the first token evaluation. It intuitively makes sense that the probabilistic mass on a single token is sensitive to its context and is biased towards a certain token (’A’) which leads to selection bias (Zheng et al., 2023). However, it is unknown whether the text answers given by the instruction-tuned models are also sensitive to such changes. Therefore, we answer this question by comparing the robustness of the text answers with the answer estimated by original/debiased first token probability (referred to first token in the following).
3.1 Experimental setup
Models
Our study focuses on open-source instruction-tuned language models since we are interested in the text response to the user’s instruction, as well as committing to open science and reproducibility. We cover three model families including Llama2 (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024). We do greedy decoding during the inference time. We also trained free and robust MCQ evaluators using Mistral-7b models which we plan to make public. For simplicity, we drop the postfix ”Inst/Chat” later.
Benchmark
We conduct the experiments on MMLU (Hendrycks et al., 2020) which consists of 4-option MCQs with tasks covering 57 subjects ranging from elementary mathematics to US history. Note that the original implementation of the debiasing method PriDe was done on the aggregated dataset level. In our robustness evaluation experiments, we believe it is instructive to go beyond the aggregate level: we employ and compare with PriDe on each task individually in order to gain deeper insight into the behavior of the model. We additionally use OpinionQA (Santurkar et al., 2023) inspired by Wang et al. (2024) to test the robustness of our trained classifier which will be introduced later.
Prompting
To give a fair comparison between the text answer and the first token evaluation, we specifically guide the model to respond directly with the option ID. We adopt the system prompt used in Wang et al. (2024) where the model is specifically asked to answer with a “single letter”: Please read the multiple-choice question below carefully and select ONE of the listed options and only give a single letter. Our preliminary experiments show that it is important to add the constraint to respond with a single letter instead of just asking it to “choose the option” used in prior works. Without such a constraint, the model will prefer to repeat the option content instead of the option ID.
Probability-based evalution
We extract the log probabilities assigned to the token corresponding to the option IDs (“A”, “B”, “C”, “D”) from the model and take the one with the highest probability as the answer. In our experiment, we see issues of only looking at the first token since different models have diverse response patterns. For example, the Llama2 models will always start the response with a space token “ ”, and the Gemma model starts its answer by repeating the word “Answer:” (e.g. “Answer: C”) most of the time. Therefore, we also use the second token of the Llama2 model and the third token of the Gemma model, which indeed leads to slightly higher robustness results, but still lower than the text answer.
Text-based evaluation
To extract the choice from the text answer automatically, we follow Wang et al. (2024) to construct a classifier by training a language model. We test five different language models: T5-small, T5-base, T5-large, Mistral-7b-Base-v0.1 and Mistral-7b-Instruct-v0.2. We finetune the Mistral models using QLoRA (Dettmers et al., 2024) implemented by Huggingface (Mangrulkar et al., 2022). See Appendix A.1 for training hyperparameters.
Table 2: Example of manually annotated data used for training our classifier.
Input Sure! The least common multiple (LCM) of 4 and 10 is 40, so the answer is (C) 40.
Acknowledgments
XW, CH and BP are supported by ERC Consolidator Grant DIALECT 101043235 and in parts by Independent Research Fund Denmark (DFF) Sapere Aude grant 9063-00077B. BM is supported by BERD@NFDI (German Research Foundation grant 460037581). PR is a member of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and is supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA).
Appendix A Appendix
A.1 Hyperparameter for the text answer classifier training
We finetuned five different language models: T5-small, T5-base, T5-large, Mistral-7b-base-v0.1 and Mistral-7b-Instruct-v0.2 for the training text answer classifier. Table 7 shows the hyperparameters and their corresponding values for the T5 models. For the two mistral models, we used QLoRA (Dettmers et al., 2024) and the default parameter settings from the Huggingface PEFT repository (Mangrulkar et al., 2022). Table 8 shows the hyperparameters and their corresponding values for the Mistral models.
Hyperparameter Value
learning_rate 2e-5
train_batch_size 1
weight_decay 0.01
bf16 True
num_train_epochs 8
Table 7: Hyperparameters for training T5 models
Hyperparameter Value
lora_r 64
lora_alpha 16
lora_dropout 0.1
task_type ”CAUSAL_LM”
use_4bit True
bnb_4bit_compute_dtype ”float16”
bnb_4bit_quant_type ”nf4”
use_nested_quant False
num_train_epochs 8
train_batch_size 4
gradient_accumulation_steps 1
gradient_checkpointing True
max_grad_norm 0.3
learning_rate 2e-4
weight_decay 0.001
optim ”paged_adamw_32bit”
lr_scheduler_type ”constant”
warmup_ratio 0.03
group_by_length True
Table 8: Hyperparameters for training the Mistral models
A.2 Annotation scheme
The annotation process was carried out by two in-house annotators, who were presented with the model outputs and the relevant multiple-choice questions. By reading the model response, the annotator has to decide which option the model is referring to. In cases where the model contradicts itself, such as choosing ‘A’ followed by describing the option content of ’B’, we annotate this case as a failure mode ’NaN’, which is not considered in this work. After the independent annotation process, two annotators discuss the samples where disagreement occurs and resolve the conflicts. Table 9 shows the data statistics of our annotated data. It is noteworthy that in the original options setting, we never observe cases where the model expresses “I do not know”, leading to 0 number of Z. However, when adding it into the options, we do often observe that model choosing “I do not know”, especially in Mistral-7b-Inst-v0.2. The option position in the Extra Options setting is shuffled.
Model Original Options Extra Options Llama2-7b-Chat A:168, B:163, C:183, D:171, Y:99, NaN:16 A:126, B:152, C:151,,D:122, E:102, F:73, G:58, NaN:16 Mistral-7b-Inst-v0.2 A:223, B:131, C:248, D:141, X:20, NaN:37 A: 156, B:96, C: 134, D:78, E:87, F: 118, G:106, NaN:25 Gemma-7b-Inst A:112, B:75, C:55, D:57, NaN:1 A:88, B:52, C:23, D:30, E: 32, F: 38, G: 36, NaN:1
Table 9: Data statistics of our annotated data.
A.3 Accuracy and selection bias in all subcategories
The detailed results the accuracy scores and selection bias of responses from all models in all subcategories are shown in the following Figures.
Figure 6: Selection bias on responses from Llama2-7b-Chat in all subcategories.
Figure 7: Accuracy on responses from Llama2-7b-Chat in all subcategories.
Figure 8: Selection bias on responses from Llama2-13b-Chat in all subcategories.
Figure 9: Accuracy on responses from Llama2-13b-Chat in all subcategories.
Figure 10: Selection bias on responses from Mistral-7b-Inst-0.2 in all subcategories.
Figure 11: Accuracy on responses from Mistral-7b-Inst-0.2 in all subcategories.
Figure 12: Selection bias on responses from Gemma-7b-Inst in all subcategories.
Figure 13: Accuracy on responses from Gemma-7b-Inst in all subcategories.